OpenStudy (anonymous):

Identify the ordered pair that represents the vector from A(-8,-1) to B(-5,3) and the magnitude of vector AB. Someone teach me how to do this problem.

OpenStudy (anonymous):

how do you find magnitude?

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

In order to find the vector that points from A to B we need to subtract each component of A from the corresponding component of B, according to the formula: \[v(a \to b) = (b _{1}-a _{1},b _{2}-a _{2)}\] In this case we have : \[v(a \to b) = (-5-(-8),3-(-1))\] \[v(a \to b) = (3,4)\] We can also see this graphicaly:  To find the magnitude we use the formula: \[||v| = \sqrt{(v _{1}^{2})+(v _{1}^{2})} \] In this case: \[||v| = \sqrt{(3^{2})+(4^{2})} \] \[||v| = \sqrt{9+16} \] \[||v| = \sqrt{25} \] \[||v| =5 \] Hope that helps.

OpenStudy (anonymous):

Ahhh sorry: All those \[||v| \] should be \[||v||\] I told you it was weird to type.

OpenStudy (anonymous):

lol its alright. thank you so much!!!! :) that helps a lot :)

OpenStudy (anonymous):

how do you find v(a)?

OpenStudy (anonymous):

One other hint: If you draw it really carefully and use real units (cm, inches or whatever) you can measure the length of your final vector right off the drawing to check the magnitude. Just keep it mind it might be a little off since it's pretty hard to get the drawing exactly perfect, but it will let you know if you're close. It works best if you can use graph paper.

OpenStudy (anonymous):

v(a), as in the position vector of point A? Or something else?

OpenStudy (anonymous):

alright. thank you :)

OpenStudy (anonymous):

you are amazing. and... im not really sure what i meant... i gotta check to see what i meant

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Ok, post a message when you find out and I'll help you with that part.

OpenStudy (anonymous):

okie dokey thank you